# Galactic transient sources with the Cherenkov Telescope Array

### v1

@@ 4.5 Novae
The only nova that has been detected at TeV energies so far is the symbiotic system RS Oph (Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022; Abe et al., 2023) which shows recurrent outburst every 15-20 years and harbours a white dwarf accreting from a late-type giant companion star. However, it could be argued that the detection is due to selection-effects based on the fact that RS Oph is relatively nearby (see below). By number, novae that involve a small, low-mass donor such as a main sequence star (these are usually the classical novae, see Chomiuk et al., 2021) are by far the most common type of system. The majority of novae have been observed in outburst only once in a human life timescale, and so far only a handful of novae are known to erupt with a recurrence time of tens of years. It has been predicted from binary population synthesis studies (Kemp et al., 2021, 2022) that most commonly, novae with evolved donors are more likely to contribute to the total current Galactic nova rate, even though by number these systems make up a smaller fraction of nova binaries.
@@
Kemp et al. (2022) estimated the Galactic nova rate to be per year. That study showed that the most common type of nova in our Galaxy today is expected to originate from a binary system involving a giant-like donor (see Fig. 11 in the aforementioned paper as a guide). As mentioned previously, currently the only system to have clearly been detected at VHE (and detected also at HE) is the symbiotic recurrent nova RS Oph. Though some groups have investigated detailed modelling of shock generation in nova systems (Hachisu & Kato, 2022; Metzger et al., 2016), which is believed to be mainly hadronic in nature, it is still not clear how many Galactic novae would be detectable by CTA at and beyond the TeV energy range. The majority of novae thus far detected at GeV energies with no clear evidence for a TeV component have been classical novae (not symbiotic systems like RS Oph, see e.g. Zheng et al., 2022). Nonetheless, we anticipate several more novae could be observed with CTA (see Chomiuk et al., 2021), particularly if these novae are detected at other wavelengths early on enabling rapid triggering and follow-up. However, assuming we can expect of the order of 30 Galactic novae per year, even with adequate triggering, it is unlikely that all of these outbursts will be detectable by CTA. If CTA-N would have been operational since August 2008 to April 2023, it would have been able to perform observations of 7 novae detected in HE gamma-ray by Fermi-LAT, in the 5 nights after their detection in optical. Assuming a similar rate of novae detected at HE gamma-ray in the future, it means that CTA-N would observe novae per year triggered by their HE gamma-ray emission.
CTA observations will be important to put constraints on the maximum energies attainable in nova explosions and the physical mechanisms involved in the production of VHE gamma rays. We estimated the capability of CTA to detect nova outbursts based on both theoretical modelling and empirical results. First, simple theoretical considerations based on the RS Oph detection are adopted to assess the gamma-ray emission at different outburst stages, following the approach in Acciari et al. (2022); H. E. S. S. Collaboration et al. (2022). Second, a parametric study based on phenomenological parameters involved in the emission of gamma rays in nova outbursts is performed to estimate the parameter space we could constrain with CTA observations. Finally, we considered dedicated numerical simulations of RS Oph to assess the expected detectability with CTA.
@@ 4.5.1 Modelling approach
We explored the capability of CTA to constrain the physical parameters of nova phenomena of different types, building up from basic arguments. The expected gamma-ray emission is obtained for different properties of a shock expanding with velocity , generated by ejected material of total mass slamming into the companion star’s wind and producing gamma rays through hadronic interactions. To accelerate protons to high energies via diffuse shock acceleration (DSA), the magnetic field has to be amplified in the shock. The maximum energy particles attain at a shock is limited ultimately by the Hillas condition (Hillas, 1984). However a more constraining limit is determined by either the time taken before radiative cooling dominates over acceleration, or by the necessary escape of the particles up-stream of the shock, in order to excite magnetic field fluctuations to a sufficient level ahead of the shock (Bell, 2004). The maximum energy in the particle spectrum, defined as as power-law function with an exponential cutoff, can be then described as:
@@
In Fig. 19 we show the expected gamma-ray flux at different times from the nova explosion, considering several physical parameters in the shock. The upper panels show the expected emission for a explosion similar to RS Oph, with , equal to 1 G and , respectively, located at a distance of 2 kpc (left) and of 4 kpc (right). The effect of increasing the star surface magnetic field (10 G) and decreasing the ejecta mass () is shown in the bottom panels, for a fixed distance of 2 kpc. For reference, the isoflux line at TeV cm-2 s-1 is marked when possible with a white line. The region below such line should be easily accessible to CTA.
@@ 4.5.2 Parametric space study
The emission detectability was assessed for both arrays of CTA using the official IRFs from prod5-v0.1
@@
for observation time) The results of the simulations for CTA-N and CTA-S are shown in panels a and b of Fig. 20, respectively. The total proton energy above 100 GeV () multiplied by , hereafter “effective proton energy reservoir”, was used as a function of to display the ratio between the integral source flux and the CTA sensitivity. This ratio was computed to obtain a qualitative estimation of the detectability of CTA for each model in the parameter space. The higher the integral flux-to-CTA-sensitivity ratio, the more feasible the detection. Moreover, the region where we would detect each model with CTA in at least one energy bin is lower-delimited in Fig. 20 by a dashed orange line to have a more precise boundary of the detection region. Therefore, the region between the dashed orange line and the white region (integral flux-to-CTA-sensitivity close to 0) delimits the border of the parameter space where CTA will likely begin to detect the gamma-ray emission of the models. Qualitatively, RS Oph would be located approximately in the top right corner of Fig. 20, while V959 Mon (the first classical nova discovered by Fermi-LAT; Ackermann et al., 2014) would be in the lower left region of the plots.
@@
The integral gamma-ray emission and the integral flux-to-CTA-sensitivity ratio in Fig. 20 increases as the effective proton energy reservoir and increase. Both top regions of plots a and b in Fig. 20 have positive values of integral flux-to-CTA-sensitivity ratio (about 30% of the total combinations), while the bottom region do not (about 70% of the total combinations). When comparing the results between CTA-N and CTA-S, the former extends the parameter space region with positive integral flux-to-CTA-sensitivity towards models with . On the other hand, the latter presents a wider detection region towards than CTA-N. CTA-N overperforms CTA-S with about 10% more detections. The better performance of CTA-N at low energies is expected because the parameter space under study was restricted to produce most of the gamma-ray emission below and it is also connected with the presence of four LSTs in the CTA-N Alpha configuration, as it is observed from current novae detected at gamma rays. Therefore, the lack of LSTs, which dominates the CTA sensitivity at these energies (Cherenkov Telescope Array Consortium et al., 2019), in the Southern array (Alpha configuration) will reduce the parameter space of detectability with CTA-S.
@@
Thus, CTA is expected to give strong constraints only to a sub-space of the whole parameter space under study. For about 30% of the area of the parameter space covered in Fig. 20 could be likely detected with CTA, in particular, where the relativistic protons have a high value of prefactor and cutoff energy. Assuming that the target proton number density is the number density of the main ejection of matter in the outburst, the results suggest that for denser ejecta, the detection region with CTA will cover a wide range of parameter values of the relativistic proton energy distribution. CTA-N should outperform CTA-S for novae with , while for , CTA-S should perform better than CTA-N at high energies.
@@ 4.5.3 RS Oph
RS Oph is a symbiotic nova formed by a high-mass white dwarf () and a red giant star (M0 III, Anupama & Mikołajewska, 1999), which transfers material to the compact object. In the literature, its distance has been estimated ranging from to (Barry et al. 2008, see also the discussion about the distance estimation in Section C.1 of the supplementary material of Acciari et al. 2022), being the most recent value of about from Gaia DR3 catalog (Gaia Collaboration, 2020) RS Oph undergoes recurrent nova outbursts with a periodicity of about 15 years. Its last eruption occurred in August 2021. Covered widely at different wavelengths, the 2021 outburst was detected at VHE gamma rays, adding a new object class to the list of VHE emitters. The HE and VHE gamma-ray emission was consistent mainly with a hadronic origin (dominated by decay), likely originated by the interaction of the ejected material with the dense wind of the red giant (Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022). The gamma-ray spectrum showed hints of hardening with time produced by the migration of gamma rays to higher energies (Acciari et al., 2022). The HE light curve presented a power-law decay after reaching the maximum emission phase. The index of the temporal decay at HE with Fermi-LAT and the one obtained at VHE by the H.E.S.S. Collaboration were compatible within errors with values and , respectively (Cheung et al., 2022; H. E. S. S. Collaboration et al., 2022). It is expected that RS Oph will undergo another outburst when CTA will be in operation. Hence, we carried out numerical simulations of RS Oph to estimate its detectability with CTA along the temporal evolution of the outburst.
4.5.4 RS Oph: CTA simulations
We performed the numerical simulations of RS Oph with CTA using the official IRFs from prod5-v0.1
for the CTA northern and southern arrays In particular, the closest IRFs set to the culmination of RS Oph in the CTA-N and CTA-S site were used (North-40deg-SouthAz
@@ 4.5.4 RS Oph: CTA simulations
) for observation time. A total of 59 daily observations of 1 hour each were simulated starting one day after the beginning of the nova outburst (batches of 100 simulations per day). We simulated this source based on the gamma-ray spectral and temporal profile reported by the MAGIC and H.E.S.S. Collaborations, respectively. The best daily-fit spectral log-parabola models from Acciari et al. (2022) were considered to model the gamma-ray emission. Spectral variations were only contemplated for the simulations of the first four days, when spectral information in Acciari et al. (2022) was available during the outburst. The spectral parameter values utilised in the different log-parabola models are shown in Table 9. After the fourth day, the spectral profile was fixed to the one from the last day with spectral information (fourth day), and the simulated gamma-ray emission was scaled to follow the power-law temporal decay reported by H.E.S.S. We set the index value of the power-law decay to .
@@
The statistical detection significance as a function of time is shown in Fig. 21. The results confirm that RS Oph would be clearly detected with CTA-N and CTA-S for the first days, reaching a detection significance of about and in an hour with CTA-N and CTA-S, respectively. RS Oph is not only detectable with CTA during the first days after the outburst, but CTA would also daily detect RS Oph up to and days after the outburst with the northern and southern arrays, respectively. If we consider the combined data of 5 and 10 adjacent days with CTA-N, the detection would be possible even up to and days, respectively. The detection would be limited down to and days with CTA-S. The resulting SED for the first simulated observation (day 1 after the outburst) is shown in Fig. 22 together with the observed spectrum obtained with MAGIC for the same observation time. The results suggest that CTA will be able to probe the gamma-ray emission for several weeks after the outburst with a precise spectral coverage at least during the first days. Using 1-h observation, CTA would be able to characterize the curvature of the VHE gamma-ray emission of RS Oph. For example, for the simulated observation of day 1, a log-parabola spectral model is preferred over a power-law model at 3.7.
Consequently, if one assumes that the next RS Oph outburst follows the same behaviour as the 2021 eruption, a plausible assumption based on the similarities observed at radio, optical and X-ray for the first weeks between 2006 and 2021 outbursts (Munari et al., 2022; Acciari et al., 2022; Page et al., 2022), CTA observations can provide detailed coverage of the gamma-ray emission during the future RS Oph outburst. Also, we could probe the maximum energy of the accelerated particles and the nova physical conditions across different outburst stages.
@@ 4.5.5 Other novae
To date, RS Oph is the only recurrent nova system from which gamma-ray emission at TeV energies has been detected during outburst. Nevertheless, several other recurrent novae, in particular symbiotic binary systems with high mass-transfer rates and dense winds, are also promising potential gamma-ray emitters. T Coronae Borealis (T CrB) in particular is a nearby symbiotic binary system, located closer to Earth than RS Oph, from which two prior outbursts have been observed in optical wavelengths (Schaefer, 2023). Models predict that the next outburst will occur in 2024.4 ± 0.3 years (Schaefer et al., 2023) and, if a shock evolution comparable to RS Oph can be assumed, particle acceleration and detectable VHE gamma-ray emission is highly expected.
@@ 4.6 Magnetars: discussion
On April 15, 2020, the Fermi-GBM and the Fermi-LAT instruments detected MeV and GeV gamma-ray emission from a giant flare event of a magnetar located in the NGC 253 galaxy (Roberts et al., 2021; Fermi-LAT Collaboration et al., 2021). The first time detection of GeV emission from a giant magnetar flare is particularly interesting, with the detection of two photons with energies 1.3 GeV and 1.7 GeV. According to Fermi-LAT Collaboration et al. (2021), these two GeV photons are produced via synchrotron emission considering the presence of a strong magnetic field which is generated in the shocks. It is proposed that these GeV photons are produced in the dissipation associated with the collision of the giant flare outflow and the external shell generated from swept-up material. This indicates that non-thermal processes accelerating particles at high-energies are at work. IC scattering can also occur in these events, and giant magnetar flares have been proposed as potential GeV-TeV emitters. TeV emission on millisecond timescales could be produced during giant flare events, which might be luminous enough to be detectable by IACTs (Lyubarsky, 2014; Murase et al., 2016), including CTA. This TeV emission could be produced via synchrotron maser mechanism, triggered by strong magnetic disturbances from the magnetosphere and propagating outwards, until they dissipate by interacting with the ambient nebula. In the case of dissipation of disturbances within the magnetar wind, a non-thermal tail can plausibly arise, potentially leading to the production of VHE gamma-ray emission (Metzger et al., 2020).
@@
The new radio facilities that will operate at the time of CTA will provide the detection of up to hundreds of FRB per day. Many of these will have good localizations and will be inside the CTA field of view, making it possible to search for prompt and/or delayed very high-energy emission. Although no magnetar outburst has been yet detected at TeV, the existence of MeV and GeV emission maintains the expectations of a possible TeV component, making magnetars good source candidates for CTA. The CTA Observatory should aim at observing magnetar flares as soon as possible, triggering on external alerts. Automatic re-pointing of the telescopes can take place whenever certain observational criteria (such as i.e. flare type, brightness, multi-wavelength counterparts or distance) are fulfilled. The high sensitivity of CTA at short timescales while provide new insights onto the physics of magnetars at VHE.

### v2

@@ 4.5 Novae
The only nova that has been detected at TeV energies so far is the symbiotic system RS Oph (Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022; Abe et al., 2025) which shows recurrent outburst every 15-20 years and harbours a white dwarf accreting from a late-type giant companion star. However, it could be argued that the detection is due to selection-effects based on the fact that RS Oph is relatively nearby (see below). By number, novae that involve a small, low-mass donor such as a main sequence star (these are usually the classical novae, see Chomiuk et al., 2021) are by far the most common type of system. The majority of novae have been observed in outburst only once in a human life timescale, and so far only a handful of novae are known to erupt with a recurrence time of tens of years. It has been predicted from binary population synthesis studies (Kemp et al., 2021, 2022) that most commonly, novae with evolved donors are more likely to contribute to the total current Galactic nova rate, even though by number these systems make up a smaller fraction of nova binaries.
@@
Kemp et al. (2022) estimated the Galactic nova rate to be per year. That study showed that the most common type of nova in our Galaxy today is expected to originate from a binary system involving a giant-like donor (see Fig. 11 in the aforementioned paper as a guide). As mentioned previously, currently the only system to have clearly been detected at VHE (and detected also at HE) is the symbiotic recurrent nova RS Oph. Though some groups have investigated detailed modelling of shock generation in nova systems (Hachisu & Kato, 2022; Metzger et al., 2016), which is believed to be mainly hadronic in nature, it is still not clear how many Galactic novae would be detectable by CTAO at and beyond the TeV energy range. The majority of novae thus far detected at GeV energies with no clear evidence for a TeV component have been classical novae (not symbiotic systems like RS Oph, see e.g. Zheng et al., 2022). Nonetheless, we anticipate several more novae could be observed with CTAO (see Chomiuk et al., 2021), particularly if these novae are detected at other wavelengths early on enabling rapid triggering and follow-up. However, assuming we can expect of the order of 30 Galactic novae per year, even with adequate triggering, it is unlikely that all of these outbursts will be detectable by CTAO. If CTAO-N would have been operational since August 2008 to April 2023, it would have been able to perform observations of 7 novae detected in HE gamma-ray by Fermi-LAT, in the 5 nights after their detection in optical. Assuming a similar rate of novae detected at HE gamma-ray in the future, it means that CTAO-N would observe novae per year triggered by their HE gamma-ray emission.
CTAO observations will be important to put constraints on the maximum energies attainable in nova explosions and the physical mechanisms involved in the production of VHE gamma rays. We estimated the capability of CTAO to detect nova outbursts based on both theoretical modelling and empirical results. First, simple theoretical considerations based on the RS Oph detection are adopted to assess the gamma-ray emission at different outburst stages, following the approach in Acciari et al. (2022); H. E. S. S. Collaboration et al. (2022). Second, a parametric study based on phenomenological parameters involved in the emission of gamma rays in nova outbursts is performed to estimate the parameter space we could constrain with CTAO observations. Finally, we considered dedicated numerical simulations of RS Oph to assess the expected detectability with CTAO.
@@ 4.5.1 Modelling approach
We explored the capability of CTAO to constrain the physical parameters of nova phenomena of different types, building up from basic arguments. The expected gamma-ray emission is obtained for different properties of a shock expanding with velocity , generated by ejected material of total mass slamming into the companion star’s wind and producing gamma rays through hadronic interactions. To accelerate protons to high energies via diffusive shock acceleration (DSA), the magnetic field has to be amplified in the shock. The maximum energy particles attain at a shock is limited ultimately by the Hillas condition (Hillas, 1984). However a more constraining limit is determined by either the time taken before radiative cooling dominates over acceleration, or by the necessary escape of the particles up-stream of the shock, in order to excite magnetic field fluctuations to a sufficient level ahead of the shock (Bell, 2004). The maximum energy in the particle spectrum, defined as as power-law function with an exponential cutoff, can be then described as:
@@
In Fig. 23 we show the expected gamma-ray flux at different times from the nova explosion, considering several physical parameters in the shock. The upper panels show the expected emission for a explosion similar to RS Oph, with , equal to 1 G and , respectively, located at a distance of 2 kpc (left) and of 4 kpc (right). The effect of increasing the star surface magnetic field (10 G) and decreasing the ejecta mass () is shown in the bottom panels, for a fixed distance of 2 kpc. For reference, the isoflux line at TeV cm-2 s-1 is marked when possible with a white line. The region below such line should be easily accessible to CTAO.
@@ 4.5.2 Parametric space study
The emission detectability was assessed for both arrays of CTAO using the official IRFs from prod5-v0.1
@@
for observation time) The results of the simulations for CTAO-N and CTAO-S are shown in panels a and b of Fig. 24, respectively. The total proton energy above 100 GeV () multiplied by , hereafter “effective proton energy reservoir”, was used as a function of to display the ratio between the integral source flux and the CTAO sensitivity. This ratio was computed to obtain a qualitative estimation of the detectability of CTAO for each model in the parameter space. The higher the integral flux-to-CTAO-sensitivity ratio, the more feasible the detection. Moreover, the region where we would detect each model with CTAO in at least one energy bin is lower-delimited in Fig. 24 by a dashed orange line to have a more precise boundary of the detection region. Therefore, the region between the dashed orange line and the white region (integral flux-to-CTAO-sensitivity close to 0) delimits the border of the parameter space where CTAO will likely begin to detect the gamma-ray emission of the models. Qualitatively, RS Oph would be located approximately in the top right corner of Fig. 24, while V959 Mon (the first classical nova discovered by Fermi-LAT; Ackermann et al., 2014) would be in the lower left region of the plots.
@@
The integral gamma-ray emission and the integral flux-to-CTAO-sensitivity ratio in Fig. 24 increases as the effective proton energy reservoir and increase. Both top regions of plots a and b in Fig. 24 have positive values of integral flux-to-CTAO-sensitivity ratio (about 30% of the total combinations), while the bottom region do not (about 70% of the total combinations). When comparing the results between CTAO-N and CTAO-S, the former extends the parameter space region with positive integral flux-to-CTAO-sensitivity towards models with . On the other hand, the latter presents a wider detection region towards than CTAO-N. CTAO-N overperforms CTAO-S with about 10% more detections. The better performance of CTAO-N at low energies is expected because the parameter space under study was restricted to produce most of the gamma-ray emission below and it is also connected with the presence of four LSTs in the CTAO-N Alpha configuration, as it is observed from current novae detected at gamma rays. Therefore, the lack of LSTs, which dominates the CTAO sensitivity at these energies (Cherenkov Telescope Array Consortium et al., 2019), in the Southern array (Alpha configuration) will reduce the parameter space of detectability with CTAO-S.
@@
Thus, CTAO is expected to give strong constraints only to a sub-space of the whole parameter space under study. For about 30% of the area of the parameter space covered in Fig. 24 could be likely detected with CTAO, in particular, where the relativistic protons have a high value of prefactor and cutoff energy. Assuming that the target proton number density is the number density of the main ejection of matter in the outburst, the results suggest that for denser ejecta, the detection region with CTAO will cover a wide range of parameter values of the relativistic proton energy distribution. CTAO-N should outperform CTAO-S for novae with , while for , CTAO-S should perform better than CTAO-N at high energies.
@@ 4.5.3 RS Oph
RS Oph is a symbiotic nova formed by a high-mass white dwarf () and a red giant star (M0 III, Anupama & Mikołajewska, 1999), which transfers material to the compact object. In the literature, its distance has been estimated ranging from to (Barry et al. 2008, see also the discussion about the distance estimation in Section C.1 of the supplementary material of Acciari et al. 2022), being the most recent value of about from Gaia DR3 catalog (Gaia Collaboration, 2020) RS Oph undergoes recurrent nova outbursts with a periodicity of about 15 years. Its last eruption occurred in August 2021. Covered widely at different wavelengths, the 2021 outburst was detected at VHE gamma rays, adding a new object class to the list of VHE emitters. The HE and VHE gamma-ray emission was consistent mainly with a hadronic origin (dominated by decay), likely originated by the interaction of the ejected material with the dense wind of the red giant (Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022). The gamma-ray spectrum showed hints of hardening with time produced by the migration of gamma rays to higher energies (Acciari et al., 2022). The HE light curve presented a power-law decay after reaching the maximum emission phase. The index of the temporal decay at HE with Fermi-LAT and the one obtained at VHE by the H.E.S.S. Collaboration were compatible within errors with values and , respectively (Cheung et al., 2022; H. E. S. S. Collaboration et al., 2022). It is expected that RS Oph will undergo another outburst when CTAO will be in operation. Hence, we carried out numerical simulations of RS Oph to estimate its detectability with CTAO along the temporal evolution of the outburst.
4.5.4 RS Oph: CTAO simulations
We performed the numerical simulations of RS Oph with CTAO using the official IRFs from prod5-v0.1
for the CTAO northern and southern arrays In particular, the closest IRFs set to the culmination of RS Oph in the CTAO-N and CTAO-S site were used (North-40deg-SouthAz
@@ 4.5.4 RS Oph: CTAO simulations
) for observation time. A total of 59 daily observations of 1 hour each were simulated starting one day after the beginning of the nova outburst (batches of 100 simulations per day). We simulated this source based on the gamma-ray spectral and temporal profile reported by the MAGIC and H.E.S.S. Collaborations, respectively. The best daily-fit spectral log-parabola models from Acciari et al. (2022) were considered to model the gamma-ray emission. Spectral variations were only contemplated for the simulations of the first four days, when spectral information in Acciari et al. (2022) was available during the outburst. The spectral parameter values utilised in the different log-parabola models are shown in Table 8. After the fourth day, the spectral profile was fixed to the one from the last day with spectral information (fourth day), and the simulated gamma-ray emission was scaled to follow the power-law temporal decay reported by H.E.S.S. We set the index value of the power-law decay to =.
@@
The statistical detection significance as a function of time is shown in Fig. 25. The results confirm that RS Oph would be clearly detected with CTAO-N and CTAO-S for the first days, reaching a detection significance of about and in an hour with CTAO-N and CTAO-S, respectively. RS Oph is not only detectable with CTAO during the first days after the outburst, but CTAO would also daily detect RS Oph up to and days after the outburst with the northern and southern arrays, respectively. If we consider the combined data of 5 and 10 adjacent days with CTAO-N, the detection would be possible even up to and days, respectively. The detection would be limited down to and days with CTAO-S. The resulting SED for the first simulated observation (day 1 after the outburst) is shown in Fig. 26 together with the observed spectrum obtained with MAGIC for the same observation time. The results suggest that CTAO will be able to probe the gamma-ray emission for several weeks after the outburst with a precise spectral coverage at least during the first days. Using 1-h observation, CTAO would be able to characterize the curvature of the VHE gamma-ray emission of RS Oph. For example, for the simulated observation of day 1, a log-parabola spectral model is preferred over a power-law model at 3.7.
Consequently, if one assumes that the next RS Oph outburst follows the same behaviour as the 2021 eruption, a plausible assumption based on the similarities observed at radio, optical and X-ray for the first weeks between 2006 and 2021 outbursts (Munari et al., 2022; Acciari et al., 2022; Page et al., 2022), CTAO observations can provide detailed coverage of the gamma-ray emission during the future RS Oph outburst. Also, we could probe the maximum energy of the accelerated particles and the nova physical conditions across different outburst stages.
@@ 4.5.5 Other novae
To date, RS Oph is the only recurrent nova system from which gamma-ray emission at TeV energies has been detected during outburst. Nevertheless, several other recurrent novae, in particular symbiotic binary systems with high mass-transfer rates and dense winds, are also promising potential gamma-ray emitters. T Coronae Borealis (T CrB) in particular is a nearby symbiotic binary system, located closer to Earth than RS Oph, from which two prior outbursts have been observed in optical wavelengths (Schaefer, 2023). Models predict that the next outburst will occur in 2024.4 ± 0.3 years (Schaefer et al., 2023) and, if a shock evolution comparable to RS Oph can be assumed, particle acceleration and detectable VHE gamma-ray emission is highly expected. Other novae that are expected to plausibly occur over the lifetime of CTAO are the recurrent novae V394 Coronae Australis, CI Aquilae or V3890 Sgr and possible superflares arising from V2487 Oph (Schaefer, 2019; Schaefer et al., 2022).
@@ 4.6 Magnetars: discussion
On April 15, 2020, the Fermi-GBM and the Fermi-LAT instruments detected MeV and GeV gamma-ray emission from a giant flare event of a magnetar located in the NGC 253 galaxy (Roberts et al., 2021; Fermi-LAT Collaboration et al., 2021). The first time detection of GeV emission from a giant magnetar flare is particularly interesting, with the detection of two photons with energies 1.3 GeV and 1.7 GeV. According to Fermi-LAT Collaboration et al. (2021), these two GeV photons are produced via synchrotron emission considering the presence of a strong magnetic field which is generated in the shocks. It is proposed that these GeV photons are produced in the dissipation associated with the collision of the giant flare outflow and the external shell generated from swept-up material. This indicates that non-thermal processes accelerating particles at high-energies are at work. IC scattering can also occur in these events, and giant magnetar flares have been proposed as potential GeV-TeV emitters. TeV emission on millisecond timescales could be produced during giant flare events, which might be luminous enough to be detectable by IACTs (Lyubarsky, 2014; Murase et al., 2016), including CTAO. This TeV emission could be produced via synchrotron maser mechanism, triggered by strong magnetic disturbances from the magnetosphere and propagating outwards, until they dissipate by interacting with the ambient nebula. In the case of dissipation of disturbances within the magnetar wind, a non-thermal tail can plausibly arise, potentially leading to the production of VHE gamma-ray emission (Metzger et al., 2020).
@@
The new radio facilities that will operate at the time of CTAO will provide the detection of up to hundreds of FRB per day. Many of these will have good localizations and will be inside the CTAO field of view, making it possible to search for prompt and/or delayed very high-energy emission. Although no magnetar outburst has been yet detected at TeV, the existence of MeV and GeV emission maintains the expectations of a possible TeV component, making magnetars good source candidates for CTAO. The CTAO Observatory should aim at observing magnetar flares as soon as possible, triggering on external alerts. Automatic re-pointing of the telescopes can take place whenever certain observational criteria (such as i.e. flare type, brightness, multi-wavelength counterparts or distance) are fulfilled. The high sensitivity of CTAO at short timescales while provide new insights onto the physics of magnetars at VHE.